# Quantized symbolic time series approximation

### v1

@@ Abstract
Time series are ubiquitous in numerous science and engineering domains, e.g., signal processing, bioinformatics, and astronomy. Previous work has verified the efficacy of symbolic time series representation in a variety of engineering applications due to its storage efficiency and numerosity reduction. The most recent symbolic aggregate approximation technique, ABBA, has been shown to preserve essential shape information of time series and improve downstream applications, e.g., neural network inference regarding prediction and anomaly detection in time series.
Motivated by the emergence of high-performance hardware which enables efficient computation for low bit-width representations, we present a new quantization-based ABBA symbolic approximation technique, QABBA, which exhibits improved storage efficiency while retaining the original speed and accuracy of symbolic reconstruction. We prove an upper bound for the error arising from quantization and discuss how the number of bits should be chosen to balance this with other errors.
An application of QABBA with large language models (LLMs) for time series regression is also presented, and its utility is investigated. By representing the symbolic chain of patterns on time series, QABBA not only avoids the training of embedding from scratch, but also achieves a new state-of-the-art on Monash regression dataset. The symbolic approximation to the time series offers a more efficient way to fine-tune LLMs on the time series regression task which contains various application domains. We further present a set of extensive experiments performed across various well-established datasets to demonstrate the advantages of the QABBA method for symbolic approximation.
Index Terms:
@@ I Introduction
The organization of the paper is as follows. Section II gives a brief review of symbolic representation methods. In Section III, we first formulate the ABBA method to unify the newly introduced notation and explore the potential for the use of quantization within the symbolic approximation of QABBA. Section IV offers insight into the storage efficiency of ABBA, fABBA, and QABBA. Section V showcases the empirical results of competing algorithms across various well-recognized datasets and Section VI concludes the paper.
II Related work
@@ III Methodology
Our approach, QABBA, shares a similar procedure with ABBA, but with a quantization step inside the transform, which enables reduced storage space for time series representation. The complete summary of notation for the QABBA procedure is listed in TABLE I. The procedures of symbolization (the first three steps) and inverse-symbolization (the last three steps)  [^1] For naming convenience, we define $\widehat{\text{{inc}}}=\widetilde{\text{{inc}}}$, following [11]. are shared among ABBA and QABBA. QABBA contains two additional steps, quantization and inverse-quantization, which can potentially add to the reconstruction error. Note that in the table we use the term “rounding” for the seventh step instead of “quantization” as was used in [11] to avoid confusion.
TABLE I: Summary of the QABBA procedure
@@ III-A Compression
$\sum_{i=i_{j-1}}^{i_{j}}$ $\Big{(}\,t_{i_{j-1}}+(t_{i_{j}}-t_{i_{j-1}})\cdot\frac{i-i_{j-1}}{i_{j}-i_{j-1}}-t_{i}\Big{)}^{2}$ (1)
$\leq(i_{j}-i_{j-1}-1)\cdot\texttt{tol}^{2}.$
@@ III-B Digitization
$\left(\texttt{scl}\frac{\texttt{len}_{1}}{\sigma_{\texttt{len}}},\frac{\texttt{inc}_{1}}{\sigma_{\texttt{inc}}}\right),$ $\left(\texttt{scl}\frac{\texttt{len}_{2}}{\sigma_{\texttt{len}}},\frac{\texttt{inc}_{2}}{\sigma_{\texttt{inc}}}\right),\hskip 2.0pt\ldots,\hskip 2.0pt$ (2)
$\left(\texttt{scl}\frac{\texttt{len}_{N}}{\sigma_{\texttt{len}}},\frac{\texttt{inc}_{N}}{\sigma_{\texttt{inc}}}\right).$
@@
III-C Quantization and inverse-quantization
@@ III-D Symbolic reconstruction
A more intuitive understanding of the symbolic approximation via QABBA can be obtained from Figure 1, which shows three ECG (Electrocardiogram) signals reconstructed using the QABBA (VQ & GA) symbolic approximation associated with an 8-bit quantization for length and 16-bit quantization for increment, using $\texttt{tol}=0.1$ and $\alpha=0.4$. We observe that the reconstructed data (for the first and second signals) via QABBA (GA) with $\texttt{scl}=1$ has a shift in the x-horizon. This can be fixed by using a higher scl value, e.g., $\texttt{scl}=5$, as illustrated in Figure 1. TABLE II lists the parameter settings, symbolic representation (output), as well as the number of distinct symbols (#Symbols) used for the three signals.
@@
TABLE II: Symbolic approximation of ECG signals.
@@
IV Storage efficiency
As discussed previously, the ABBA method and its variants are symbolic approximation techniques that achieve efficient data compression, significantly reducing the storage need of time series. It is known that reconstructing an ABBA representation requires the initial value(s), symbolic center(s), and strings (we refer to Section III for more detail). Therefore, to store the time series representation, only these three variables are needed.
@@ IV Storage efficiency
V Empirical Results
In this section, we compare the performance of non-quantized ABBA, i.e., ABBA and fABBA, quantized ABBA, and SAX-based algorithms, in terms of storage and approximation errors. We conduct extensive experiments on synthetic time series as well as and well-established real-world datasets including those from the UCR Archive [47] and the UEA Archive [48]. For a fair comparison, the competing algorithms are selected from those for which the software implementations are publicly available [^2] ABBA and fABBA are available at https://github.com/nla-group/ABBA and https://github.com/nla-group/fABBA..
In the following experiments, we compare QABBA with ABBA, fABBA, SAX, and $1d$-SAX algorithms. Additionally, we test two variants of QABBA—QABBA (VQ) and QABBA (GA)—which use k-means++ [49] and greedy aggregation [12], respectively, in the digitization stage. All experiments here except Section V-B are simulated in Python on a Dell PowerEdge R740 Server with Intel Xeon Silver @ 4114 2.2GHz and 1.5 TB RAM.
V-A Approximation error from quantization
@@
V-B LLM with QABBA for time series regression
In this section, we show that QABBA can exploit the semantic information hiding beneath the time series in the task of time series regression. For the regression task, we evaluate one pre-trained LLM using the Time Series Extrinsic Regression (TSER) benchmark archive [50], which contains 19 time series datasets from 5 distinct areas, ranging from health monitoring, energy monitoring, environment monitoring, sentiment analysis to forecasting [^3] Monash regression data is available at http://tseregression.org/.. Details of the hyperparameter settings of our implementation can be found in TABLE III. We simulated QABBA in various bit-width settings for len and inc and applied one commonly-used LLM - Mistral-7B [51] - to tackle time series regression problems. In the first, the numerical time series is transformed to the symbolic series, and the corresponding label is still a number. By replacing the last layer with a regression layer, we fine-tune Mistral-7B on the training data using QLoRA [52] with inhibition [53]. All the experiments are simulated in PyTorch [54] and conducted on a single NVIDIA A100 40GB GPU.
The empirical results are shown in TABLE IV, in which 9 out of 16 use-cases outperform the machine learning state-of-the-art results (SOTA). Specifically, when using 16 bit-width for both len and inc, QABBA not only shows competitive performance for the regression of time series but also occupies less memory space. Also, we can observe that lower bit-width use for len and inc only does not degrade the performance of regression much. QABBA is able to provide chain-of-patterns to LLMs by compressing and digitizing time series to symbols, which finally results in the change of embedding space by using adaption fine-tuning methods. The benefits of QABBA with LLMs include (1) avoiding the need for LLMs to learn time series from scratch, (2) only utilizing compression and decompression without the need for the training of extra embeddings [55].
TABLE III: Hyperparameters of the time series regression task. Quant. is the LoRA quantization process. Inhib. is the inhibition threshold in QLoRA. MAE is mean absolute error.
@@ V-B LLM with QABBA for time series regression
TABLE IV: The performance (RMSE: root mean square error) of time series regression task on QABBA. SOTA is the results of [50].
@@
[…]
LLM Qunatize Monash Regression Data
@@
[…]
V-C Performance profiles in UCR Archive
@@ V-D Simulations in UEA Archive
In this experiment, we evaluate the competing ABBA-based algorithms in terms of approximation error and compression ratio for selected datasets from the UEA Archive (see TABLE V). The approximation error is measured by the distance between ground truth data and reconstructed data (simply computed as the average of the squared differences given multidimensional tensor). The compression time is not measured since all algorithms perform the same compression and their differences lie only in their digitization. All algorithms use the same number of symbolic centers. We use the compression ratio defined in Section IV to evaluate the storage savings. Experiments using two compression settings ($\text{{tol}}=0.005$ and $\text{{tol}}=0.05$) are shown in Figure 7 and Figure 8.
@@
The acceptable error can be evaluated perceptually by plotting both the raw time series and the reconstruction; if they are close enough, this might be “acceptable“. Both ABBA and fABBA are well-established methods for quality symbolic reconstruction, and our intent is to validate the reconstruction error by comparing it with ABBA or fABBA. Hence we plot the reconstruction of ABBA, fABBA, QABBA (VQ), and QABBA (GA) performed with $\text{{tol}}=0.05$ in the first (others’ performance is similar through our test) multivariate time series of the UEA datasets in TABLE V following the settings described in Section V-D. The reconstruction comparisons are as shown in Figure 6. The reconstruction of ABBA and QABBA almost overlaps with the original time series, which demonstrates that all the tested methods achieve satisfying reconstruction, and the comparison of their storage is justified.
@@
Figure 6: Symbolic reconstruction of UEA Archive.
TABLE V: Selected datasets in UEA Archive.
@@
VI Summary and future work

### v2

@@ Abstract
Time series are ubiquitous in numerous science and engineering domains, e.g., signal processing, bioinformatics, and astronomy. Previous work has verified the efficacy of symbolic time series representation in a variety of engineering applications due to its storage efficiency and numerosity reduction. The most recent symbolic aggregate approximation technique, ABBA, has been shown to preserve essential shape information of time series and improve downstream applications, e.g., neural network inference regarding prediction and anomaly detection in time series. Motivated by the emergence of high-performance hardware which enables efficient computation for low bit-width representations, we present a new quantization-based ABBA symbolic approximation technique, QABBA, which exhibits improved storage efficiency while retaining the original speed and accuracy of symbolic reconstruction. We prove an upper bound for the error arising from quantization and discuss how the number of bits should be chosen to balance this with other errors. We present an application of QABBA combined with large language models (LLMs) for time series regression, exploring its effectiveness. By encoding time series as symbolic pattern chains, QABBA eliminates the need to train embeddings from scratch and achieves state-of-the-art results on the Monash regression dataset. This symbolic approximation enhances the efficiency of fine-tuning LLMs for time series regression across diverse domains. Extensive experiments on well-established datasets further highlight the benefits of QABBA’s symbolic approach.
keywords:
@@ 1 Introduction
The organization of the paper is as follows. Section 2 gives a brief review of symbolic representation methods. In Section 3, we first formulate the ABBA method to unify the newly introduced notation and explore the potential for the use of quantization within the symbolic approximation of QABBA. Section 4 offers insight into the storage efficiency of ABBA, fABBA, and QABBA. Section 5 showcases the empirical results of competing algorithms across various well-recognized datasets and Section 6 concludes the paper.
2 Related work
@@ 3 Methodology
Our approach, QABBA, shares a similar procedure with ABBA, but with a quantization step inside the transform, which enables reduced storage space for time series representation. The complete summary of notation for the QABBA procedure is listed in Table 1. The procedures of symbolization (the first three steps) and inverse-symbolization (the last three steps)  [^1] For naming convenience, we define $\widehat{\text{{inc}}}=\widetilde{\text{{inc}}}$, following [11]. are shared among ABBA and QABBA. QABBA contains two additional steps, quantization and inverse-quantization, which can potentially add to the reconstruction error. Note that in the table we use the term “rounding” for the seventh step instead of “quantization” as was used in [11] to avoid confusion.
Table 1: Summary of the QABBA procedure
@@ 3.1 Compression
$\sum_{i=i_{j-1}}^{i_{j}}$ $\Big{(}\,t_{i_{j-1}}+(t_{i_{j}}-t_{i_{j-1}})\cdot\frac{i-i_{j-1}}{i_{j}-i_{j-1}}-t_{i}\Big{)}^{2}\leq(i_{j}-i_{j-1}-1)\cdot\texttt{tol}^{2}.$ (1)
@@ 3.2 Digitization
$\left(\texttt{scl}\frac{\texttt{len}_{1}}{\sigma_{\texttt{len}}},\frac{\texttt{inc}_{1}}{\sigma_{\texttt{inc}}}\right),$ $\left(\texttt{scl}\frac{\texttt{len}_{2}}{\sigma_{\texttt{len}}},\frac{\texttt{inc}_{2}}{\sigma_{\texttt{inc}}}\right),\hskip 2.0pt\ldots,\hskip 2.0pt\left(\texttt{scl}\frac{\texttt{len}_{N}}{\sigma_{\texttt{len}}},\frac{\texttt{inc}_{N}}{\sigma_{\texttt{inc}}}\right).$ (2)
@@
3.3 Quantization and inverse-quantization
@@ 3.4 Symbolic reconstruction
A more intuitive understanding of the symbolic approximation via QABBA can be obtained from Figure 1, which shows three ECG (Electrocardiogram) signals reconstructed using the QABBA (VQ & GA) symbolic approximation associated with an 8-bit quantization for length and 16-bit quantization for increment, using $\texttt{tol}=0.1$ and $\alpha=0.4$. We observe that the reconstructed data (for the first and second signals) via QABBA (GA) with $\texttt{scl}=1$ has a shift in the x-horizon. This can be fixed by using a higher scl value, e.g., $\texttt{scl}=5$, as illustrated in Figure 1. Table 2 lists the parameter settings, symbolic representation (output), as well as the number of distinct symbols (#Symbols) used for the three signals.
@@
Table 2: Symbolic approximation of ECG signals.
@@
4 Storage efficiency
As discussed previously, the ABBA method and its variants are symbolic approximation techniques that achieve efficient data compression, significantly reducing the storage need of time series. It is known that reconstructing an ABBA representation requires the initial value(s), symbolic center(s), and strings (we refer to Section 3 for more detail). Therefore, to store the time series representation, only these three variables are needed.
@@ 4 Storage efficiency
5 Empirical Results
In this section, we compare the performance of non-quantized ABBA, i.e., ABBA and fABBA, quantized ABBA, and SAX-based algorithms, in terms of storage and approximation errors. We conduct extensive experiments on synthetic time series as well as and well-established real-world datasets including those from the UCR Archive [47] and the UEA Archive [48]. To ensure a fair comparison, we selected competing algorithms with publicly available software implementations. [^2] ABBA and fABBA are accessible at https://github.com/nla-group/ABBA and https://github.com/nla-group/fABBA, respectively. Our software is integrated into the established fABBA library (https://github.com/nla-group/fABBA), and the experimental code is openly available at:
https://github.com/inEXASCALE/qabba.
In the following experiments, we compare QABBA with ABBA, fABBA, SAX, and $1d$-SAX algorithms. Additionally, we test two variants of QABBA—QABBA (VQ) and QABBA (GA)—which use k-means++ [49] and greedy aggregation [12], respectively, in the digitization stage. All experiments here except Section 5.2 are simulated in Python on a Dell PowerEdge R740 Server with Intel Xeon Silver @ 4114 2.2GHz and 1.5 TB RAM.
5.1 Approximation error from quantization
@@
5.2 LLM with QABBA for time series regression
In this section, we show that QABBA can exploit the semantic information hiding beneath the time series in the task of time series regression. For the regression task, we evaluate one pre-trained LLM using the Time Series Extrinsic Regression (TSER) benchmark archive [50], which contains 19 time series datasets from 5 distinct areas, ranging from health monitoring, energy monitoring, environment monitoring, sentiment analysis to forecasting [^3] Monash regression data is available at http://tseregression.org/.. Details of the hyperparameter settings of our implementation can be found in Table 3. We simulated QABBA in various bit-width settings for len and inc and applied one commonly-used LLM - Mistral-7B [51] - to tackle time series regression problems. In the first, the numerical time series is transformed to the symbolic series, and the corresponding label is still a number. By replacing the last layer with a regression layer, we fine-tune Mistral-7B on the training data using QLoRA [52] with inhibition [53]. All the experiments are simulated in PyTorch [54] and conducted on a single NVIDIA A100 40GB GPU.
The empirical results are shown in Table 4, in which 9 out of 16 use-cases outperform the machine learning state-of-the-art results (SOTA). Specifically, when using 16 bit-width for both len and inc, QABBA not only shows competitive performance for the regression of time series but also occupies less memory space. Also, we can observe that lower bit-width use for len and inc only does not degrade the performance of regression much. QABBA is able to provide chain-of-patterns to LLMs by compressing and digitizing time series to symbols, which finally results in the change of embedding space by using adaption fine-tuning methods. The benefits of QABBA with LLMs include (1) avoiding the need for LLMs to learn time series from scratch, (2) only utilizing compression and decompression without the need for the training of extra embeddings [55].
Table 3: Hyperparameters of the time series regression task. Quant. is the LoRA quantization process. Inhib. is the inhibition threshold in QLoRA. MAE is mean absolute error.
@@ 5.2 LLM with QABBA for time series regression
Table 4: The performance (RMSE: root mean square error) of time series regression task on QABBA. SOTA is the results of [50].
@@
SOTA - - 2.29 132.80 32.61 0.64 93.14 59.50 28.80 0.00
LLM Quantize Monash Regression Data
@@
SOTA - - 0.01 0.00 8.12 9.92 23.90 0.14 0.14 0.04
5.3 Performance profiles in UCR Archive
@@ 5.4 Simulations in UEA Archive
In this experiment, we evaluate the competing ABBA-based algorithms in terms of approximation error and compression ratio for selected datasets from the UEA Archive (see Table 5). The approximation error is measured by the distance between ground truth data and reconstructed data (simply computed as the average of the squared differences given multidimensional tensor). The compression time is not measured since all algorithms perform the same compression and their differences lie only in their digitization. All algorithms use the same number of symbolic centers. We use the compression ratio defined in Section 4 to evaluate the storage savings. Experiments using two compression settings ($\text{{tol}}=0.005$ and $\text{{tol}}=0.05$) are shown in Figure 8 and Figure 9.
@@
The acceptable error can be evaluated perceptually by plotting both the raw time series and the reconstruction; if they are close enough, this might be “acceptable“. Both ABBA and fABBA are well-established methods for quality symbolic reconstruction, and our intent is to validate the reconstruction error by comparing it with ABBA or fABBA. Hence we plot the reconstruction of ABBA, fABBA, QABBA (VQ), and QABBA (GA) performed with $\text{{tol}}=0.05$ in the first (others’ performance is similar through our test) multivariate time series of the UEA datasets in Table 5 following the settings described in Section 5.4. The reconstruction comparisons are as shown in Figure 6 and 6. The reconstruction of ABBA and QABBA almost overlaps with the original time series, which demonstrates that all the tested methods achieve satisfying reconstruction, and the comparison of their storage is justified.
@@
Figure 7: Symbolic reconstruction of UEA Archive (II).
Table 5: Selected datasets from the UEA Archive.
@@
6 Summary and future work
@@
Acknowledgements
The first author acknowledges funding from the European Union (ERC, inEXASCALE, 101075632). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. Additionally, the first author acknowledges funding from the Charles University Research Centre program (No. UNCE/24/SCI/005). The second author acknowledges funding from the France 2030 NumPEx Exa-MA project (ANR-22-EXNU-0002), managed by the French National Research Agency (ANR).